Exercise 16.56

Consider the reaction:
CO(g)+H2O(g)⇌CO2(g)+H2(g)
Kc=102 at 500 K
A reaction mixture initially contains 0.135 M CO and 0.135 M H2O.

Part A

What will be the equilibrium concentration of [CO]?

[CO] =   M  

SubmitMy AnswersGive Up

Part B

What will be the equilibrium concentration of [H2O]?

[H2O] =   M  

SubmitMy AnswersGive Up

Part C

What will be the equilibrium concentration of [CO2]?

[CO2] =   M  

SubmitMy AnswersGive Up

Part D

What will be the equilibrium concentration of [H2]?

[H2] =   M  
Answer #1

0-135 0-135. 0.135-x 0.135-χ 0 13S- ) 0.135ース z= 0.12.3 [COJ = 0-135-0-12.3 太 [co] = 0.0120 M